OpenStudy (jadareese_18):

Compare the graphs of the logarithmic functions f(x)=log7x and g(x)=log4x. For what values of x is f=g, f>g, and f

OpenStudy (solomonzelman):

Do you agree with this? \(f(x)=\log(7x)=\log(7)+\log(x)\) \(g(x)=\log(4x)=\log(4)+\log(x)\)

OpenStudy (solomonzelman):

This is because \(\log(a\times b)=\log(a)+\log(b)\). Right?

OpenStudy (solomonzelman):

So, if you want to compare them (assuming that \(\log(x)\) is defined, that is \(x>0\)), then you have: \({\tiny~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~}f(x)~~■~~g(x)\) \(\\[0.8em]\) \({\tiny~~}\log(7x)~~■~~\log(4x)\) \(\\[0.8em]\) \(\log(7)+\log(x)~~■~~\log(4)+\log(x)\) \(\\[0.8em]\) \({\tiny~~~}\log(7)~~■~~\log(4)\) \(\\[0.8em]\) (And note that this comparison is for any value of \(x\))

OpenStudy (jadareese_18):

I think I'm following but a little confused with the a and b part.

OpenStudy (solomonzelman):

I am using \( ■\) just as an unknown symbol \(-,~+,~\le,~\ge,~=\) etc. (Similarly and for the same purpose, the symbol \( ♣\) is in use.)

OpenStudy (jadareese_18):

Also the numbers after the (log) part of the equation were suppose to be lowered. I wasn't sure how to do that.

OpenStudy (solomonzelman):

Oh, that is just a general rule, that for any positive A and B, if you have a log of a product of A and B, then this is equivalent to the sum of individual logs of A and of B. or, mathematically, \(\log(a\times b)=\log(a)+\log(b)\).

OpenStudy (solomonzelman):

What do you mean that "the (log) part of the equation were suppose to be lowered" ?

OpenStudy (jadareese_18):

the number after the log part so it should have been f(x)=log v7 x and g(x)=log v4 x where the 7 and 4 are lowered below the log and x.

OpenStudy (solomonzelman):

Oh ... I see.

OpenStudy (solomonzelman):

So they are not logs of 7x and of 4x. Rather, they are log base 7 of x, and base 4 of x.

OpenStudy (solomonzelman):

\(f(x)=\log_7x \\ g(x)=\log_4x\)

OpenStudy (jadareese_18):

Yes that is correct. I'm sorry.

OpenStudy (solomonzelman):

If you are familiar with latex/tex, then here is the code to use. `\(f(x)=\log_7x\)` (You make a subscript with \(\color{red}{\text{_}}\) or with \(\color{red}{\text{_{ }}}\).

OpenStudy (jadareese_18):

Oh thank you! That's really nice to know.

OpenStudy (solomonzelman):

\(f(x)=\log_7x \quad \Longleftrightarrow \quad7^{f(x)}=x\quad \Longleftrightarrow \quad f^{-1}(x)=7^{x}\). \(g(x)=\log_4x \quad \Longleftrightarrow \quad4^{g(x)}=x\quad \Longleftrightarrow \quad g^{-1}(x)=4^{x}\). We can compare their inverse functions, that might be easier.